OpenStudy (anonymous):

what is the 7th and 10th term in this situation an=2+(n-1)3

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

is this really \[a_n=2+(n-1)^3\] or \[a_n=2+(n-1)\times 3\]?

OpenStudy (anonymous):

that is it

OpenStudy (anonymous):

there are two different expressions which one is it?

OpenStudy (anonymous):

the 2nd one

OpenStudy (anonymous):

Since \[ a_n = 2+(n-1)\cdot 3 \]We can say the 7th term, \(a_7\) is: \[ a_7 = 2+(7-1)\cdot 3 \]

OpenStudy (anonymous):

where you see an \(n\) replace it by \(7\) and get what @wio wrote above

OpenStudy (anonymous):

But they will want you to simplify.

OpenStudy (anonymous):

wait so wio are you correct?

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

wats the 10th

OpenStudy (anonymous):

What I gave you is how to find the solution, you still have to do the math to simplify.

OpenStudy (anonymous):

The 10th term can be found with the same pattern.

OpenStudy (anonymous):

\[ a_{\#} =2+(\#-1)\cdot 3 \]

OpenStudy (anonymous):

wat do i have to do now simplify?

OpenStudy (anonymous):

For example: \[ 2+(7-1)\cdot 3 = 2+ (6)\cdot 3 \]

OpenStudy (anonymous):

thanx but this too hard